ShantiKai66:

A sports apparel company surveyed the market about the size of T-shirts people most frequently buy: small, medium, large, or extra-large. The table shows the survey results. T-Shirt Size Number of Responses small 11 medium 37 large 83 extra-large 19 To help its sales, the company should focus on producing size _______ (choices: small, medium, large, XL.) T-shirts. Approximately ______(choices: 55.3, 80, 83, 92.) % of T-shirts the company makes should be medium and large T-shirts to match the market preference. If the total number of T-shirts sold in a month is 45,000, _____(choices: 1,750, 3,300, 4,950, 14, 850.) of these will likely be size small.

ZombieLover:

Hi welcome, do you know the first step?

ShantiKai66:

Well, I know that you're supposed to add all of them up like this: 11 + 37 + 83 + 92? Right? That equals 150, but I get confused from there.

ZombieLover:

ok well yes you add those up, the first part you will go with the largest amount which will be (83) so that's 83 out of 150 people

ShantiKai66:

Yep. Which would mean that large would be the answer to the 1st multiple choice.

ZombieLover:

in order to find the percentage of the medium AND large shirts, add them together (37+83=?) and then divide by the total of 150. (120/150 =12/15=4/5 which equals ?%.)

ZombieLover:

Correct :)

ShantiKai66:

Ok

ZombieLover:

here so its easier \[37 +83=?\]

ShantiKai66:

37 + 83 = 120

ZombieLover:

and \[120/150= 12/15=4/5= ?%\]

ZombieLover:

correct

ShantiKai66:

120/150 = 0.8 or 80%. 12/15 = 0.8 or 80%. and 4/5 = 0.8 or 80%.

ZombieLover:

and that would be your answer for number of t-shirts :)

ShantiKai66:

So, second answer would be 80%, but what would the third answer be?

ZombieLover:

now to find the Third we will have the total number of sales in each month which is $45,000, and multilpy that by 11 out of 150 to get the answer

ZombieLover:

\[11/150 X 45,000 = ?\]

ShantiKai66:

3,300.

ZombieLover:

Correct :)

ShantiKai66:

Thanks SOOO much!!

ZombieLover:

Your welcome!

ShantiKai66:

:)